OpenStudy (anonymous):

why square root (3/2)can be change to square root 6/2

OpenStudy (babalooo):

hello

OpenStudy (babalooo):

good question. take a look\[\sqrt{3/2} =\sqrt{3}/\sqrt{2}\] and you rationalize the botton so no sqrt exist, you multiply both top and bottom by sqrt(2) when you do that you end up with \[(\sqrt{3} \times \sqrt{2})/2 = \sqrt{6}/2\]

OpenStudy (anonymous):

idk sorry